OpenStudy (anonymous):

how to find the inverse function of 3x+lnx

OpenStudy (anonymous):

then find f-1(3)

OpenStudy (anonymous):

to take inverse of this question you need to combine the expression such that you get a new function. in form f=ln(something)

OpenStudy (anonymous):

and how would you do that?

OpenStudy (anonymous):

ok check this out: 3x=ln(e^(3x))

OpenStudy (anonymous):

totally lost, don't know how you got it

OpenStudy (anonymous):

y=3x+lnx x= 3y+lny x=y(3+ln) y= x/ 3+ln not sure

OpenStudy (anonymous):

ok see you know this form of ln function: ln(e)=1 ln(e^5)=5ln(e)

OpenStudy (anonymous):

5ln(e)=5

OpenStudy (anonymous):

non nancy that's not possible you cannot take y common like that

OpenStudy (anonymous):

nope, sorry, still doesn't make sense

OpenStudy (anonymous):